OpenStudy (anonymous):

[4 5 ] +2 [1 8] [0 -1] [-4 -6] [-3 3] - [2 6 ] [5 9] [-5 11] -0.5[10 8] [-2 -6] can someone simplify these?

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started \[\Large \begin{bmatrix} 4 & 5\\ 0 & -1 \end{bmatrix} +2 \begin{bmatrix} 1 & 8\\ -4 & -6 \end{bmatrix}\] \[\Large \begin{bmatrix} 4 & 5\\ 0 & -1 \end{bmatrix} + \begin{bmatrix} 2*1 & 2*8\\ 2*(-4) & 2*(-6) \end{bmatrix}\] \[\Large \begin{bmatrix} 4 & 5\\ 0 & -1 \end{bmatrix} + \begin{bmatrix} 2 & 16\\ -8 & -12 \end{bmatrix}\] \[\Large \begin{bmatrix} 4+2 & 5+16\\ 0+(-8) & -1+(-12) \end{bmatrix}\] \[\Large \begin{bmatrix} 6 & 21\\ -8 & -13 \end{bmatrix}\] So \[\Large \begin{bmatrix} 4 & 5\\ 0 & -1 \end{bmatrix} +2 \begin{bmatrix} 1 & 8\\ -4 & -6 \end{bmatrix} = \begin{bmatrix} 6 & 21\\ -8 & -13 \end{bmatrix} \]